OpenStudy (anonymous):

The division property of equality could be used to solve which of the following equations? 5x= 30 x+ 3 = 7 (x+ 2) (x- 2) = 0 = 16

OpenStudy (anonymous):

5x=30 You ahve to divide both the sides by 5 to get the answer

OpenStudy (anonymous):

tht is an answer to the q all 4 of them are i just dnt know wich one

OpenStudy (anonymous):

i thiught it ws the 3rd one

OpenStudy (anonymous):

To find the answer to 3rd one, you don't have to apply division rule

OpenStudy (anonymous):

The second one depends on subtraction..the fourth choice is incomplete!

OpenStudy (anonymous):

(x+2)(x-2) = 0 This implies either x+2=0 or x-2=o Therefore, the answers are x=-2 or x=+2 See, no division in any step, but you got the answer

OpenStudy (anonymous):

it would have to be th 4th one cuz it x/4=16

OpenStudy (anonymous):

x/4 = 16 (multiplication rule) Multiply both sides with 4, x/4 * 4 = 16*4 x = 64 You don't need division here, its multiplication

OpenStudy (anonymous):

an u have to do 16/4 to gt th answer

OpenStudy (anonymous):

oh i see ok

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

ABDC is an isosceles trapezoid. Given only the choices below, which properties would you use to prove ACD ≅ BDC by SSS? The legs are congruent. The base angles are congruent. The bases are | |.The diagonals are congruent.